How do you find the roots and degree for : h(x)=x(x²-x-6)
Answer by edjones(8007)

h(x)=x(x²-x-6)
x(x^2-x-6)=0
x(x-3)(x+2)=0
x=0, x=3, x=-2
It is a 3rd degree equation because it has 3 roots.
.
Ed